A square parallel-plate capacitor 5.0 cm on a side
Answer #1

Displacement current is dened as ID = E0 * dΦE / dt

ID = E0 * A/d * dV / dt

= E0 * [(0.05)^2/(0.000450)] * (525000)

= 8.854 * 10^-12 * [(0.05)^2/(0.000450)] * (525000)

= 25.8 *10^-6 A (or 26uA)